OpenStudy (anonymous):

Will Medal! Is cos^2x(1+ tan^2x) = 1 and Identitie

hartnn (hartnn):

what identities do you know about tan^2 x ?

OpenStudy (anonymous):

Honestly nun I just know Identities are always true

hartnn (hartnn):

you can use 1+ tan^2 x = sec^2 x and sec x = 1/cos x

OpenStudy (anonymous):

and That makes it an Identitie?

hartnn (hartnn):

doesn't it ? = cos^2 x (sec^2 x) = cos^2 x * (1/cos^2 x) = 1